OpenStudy (anonymous):

Write an equation in point-slope form of the line through the given point of the given slop? 3.(5,1);m= 1/3 4.(-2,3);m= -2

OpenStudy (texaschic101):

point slope form : y - y1 = m(x - x1) (5,1) x1 = 5 and y1 = 1 slope(m) = 1/3 now we sub y - 1 = 1/3(x - 5) -- point slope form ======================= y - y1 = m(x - x1) (-2,3) x1 = -2 and y1 = 3 slope(m) = -2 now we sub y - 3 = -2(x - (-2) y - 3 = -2(x + 2) -- point slope form

OpenStudy (anonymous):

thank you! help me with more ?

OpenStudy (texaschic101):

lets see what ya got

OpenStudy (anonymous):

y=3/4x+5 y+4=1/3(x+6